OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form.

hero (hero):

y - 4 = -2(x-5) y - 4 = -2x +10 y = -2x+10+4 y = -2x+14

hero (hero):

Easy

OpenStudy (anonymous):

Explain?

OpenStudy (anonymous):

Oh, nvm

OpenStudy (anonymous):

I see it now.

hero (hero):

slope intercept form is y = mx+b